Larkspur Inc. is building a new hockey arena at a cost of $1,900,000. It received a down payment of $380,000 from local businesses to support the project, and now needs to borrow $1,520,000 to complete the project. It therefore decides to issue $1,520,000 of 10-year, 10.5% bonds. These bonds were issued on January 1, 2020, and pay interest annually on each January 1. The bonds yield 10% to the investor and have an effective interest rate to the issuer of 10.4053%. (There is an increased effective interest rate due to the capitalization of the bond issue costs.) Any additional funds that are needed to complete the project will be obtained from local businesses. Larkspur Inc. paid and capitalized $38,000 in bond issuance costs related to the bond issue. Larkspur prepares financial statements in accordance with IFRS.

A.) Using (1) factor tables, (2) a financial calculator, or (3) Excel function PV, calculate the value of the bonds and prepare the journal entry to record the issuance of the bonds on January 1, 2020. (Hint:Refer to Chapter 3 for tips on calculating. For the journal entry, use the amount arrived at using the time value of money tables.)

B.) Prepare a bond amortization schedule up to and including January 1, 2025, using the effective interest method.

C.) Assume that on July 1, 2023, the company retires half of the bonds at a cost of $809,000 plus accrued interest. Prepare the journal entries to record this retirement.
Answer #1

Part A

Present value function

Pmt = 1520000*10.50% = $159600

Nper = 10

Fv = 1520000

Rate=10%

PV function of excel

=-PV(rate,nper,PMT, (FV),type)

=-PV(10%,10,159600,1520000)

=$1566699

Date

General journal

Debit

Credit

Jan. 1

Cash

1528699

Unamortized bond issue costs

38000

Bonds payable

1520000

Premium bonds payable

46699

Part B

Date

Interest paid

Interest expense (10%)

Premium amortization

Bond carrying

Jan1, 20

1566699

Jan1, 21

159600

156670

2930

1563769

Jan1, 22

159600

156377

3223

1560546

Jan1, 23

159600

156055

3545

1557001

Jan1, 24

159600

155700

3900

1553101

Jan1, 25

159600

155310

4290

1548811

Interest = 1520000*10.50% = 159600

Interest expense = previous bond carrying value * 10%

Premium amortization = interest paid – interest expense

Bond carrying value = previous bond carrying value – premium amortization

Part C

Unamortized Bond Issue Costs =(38000- (38000/10*6/12*7))*50% = 12350

Carrying value of bonds as of 1/1/23 (1557001*50%)

778501

Amortization of bond premium until 7/1/19 (159600/4)-(155700/4)

975

Carrying value of bonds as of 7/1/19

777526

Carrying value of ½ of bonds as of 7/1/19

777526

- reacquisition price of ½ of bonds (809000+12350)

821350

= gain or loss on redemption of ½ of bonds

-43824

Entry of accrued interest

Date

General journal

Debit

Credit

Jul. 1

Interest expense (778501*10%*6/12)

38925

Premium on bonds payable

975

Cash

39900

Entry for reacquisition

Date

General journal

Debit

Credit

Jul. 1

Bonds Payable ($1520000 × 50%)

760000

Premium on bonds payable

17526

Loss on Redemption of Bonds

43824

Unamortized Bond Issue Costs

12350

Cash

809000